OpenStudy (anonymous):

Two gears are connected and are rotating simultaneously. The smaller gear has a radius of 4 inches, and the larger gear has a radius of 7 inches. two circles touching at one point. Larger circle has radius of 7 inches. Smaller circle has radius of 4 inches. Part 1: What is the angle measure, in degrees and rounded to the nearest tenth, through which the larger gear has rotated when the smaller gear has made one complete rotation? Part 2: How many rotations will the smaller gear make during one complete rotation of the larger gear? Show all work.

OpenStudy (anonymous):

Would I divide 14pi/8pi?

OpenStudy (anonymous):

So far I have arc length 25.133 for the smaller circle, and 43.982 for the bigger circle. 43.982 - 25.133 = 18.849?

OpenStudy (anonymous):

circumference of smaller gear=2 pi r=2*pi*4=8 pi